# Optimal Client Sampling in Federated Learning with Client-Level Heterogeneous Differential Privacy

### v1

@@ Abstract
Federated Learning with client-level differential privacy (DP) provides a promising framework for collaboratively training models while rigorously protecting clients’ privacy. However, classic approaches like DP-FedAvg struggle when clients have heterogeneous privacy requirements, as they must uniformly enforce the strictest privacy level across clients, leading to excessive DP noise and significant model utility degradation. Existing methods to improve the model utility in such heterogeneous privacy settings often assume a trusted server and are largely heuristic, resulting in suboptimal performance and lacking strong theoretical underpinnings. In this work, we address these challenges under a practical attack model where both clients and the server are honest-but-curious. We propose GDPFed, which partitions clients into groups based on their privacy budgets and achieves client-level DP within each group to reduce the privacy budget waste and hence improve the model utility. Based on the privacy and convergence analysis of GDPFed, we find that the magnitude of DP noise depends on both model dimensionality and the per-group client sampling ratios. To further improve the performance of GDPFed, we introduce GDPFed+, which integrates model sparsification to eliminate unnecessary noise and optimizes per-group client sampling ratios to minimize convergence error. Extensive empirical evaluations on multiple benchmark datasets demonstrate the effectiveness of GDPFed+, showing substantial performance gains compared with state-of-the-art methods.
1 Introduction
@@
Traditional centralized Machine Learning (ML) frameworks require collecting all training data at a single node (e.g., a central server), raising significant privacy concerns. To mitigate this issue, Federated Learning (FL) [37] has emerged as a distributed ML paradigm that enables model training directly on decentralized data sources without transferring raw data. In FL, multiple local clients (e.g., edge devices) collaboratively train a shared global model under the coordination of a central server. Specifically, in each training round, the server sends the global model to a subset of clients, who update it using their private data. These model updates are then transmitted to the server, which aggregates them to refine the global model. This process continues until the global model converges.
@@
Although the FL paradigm keeps sensitive training data on clients, recent studies have shown that adversaries can still infer private information through well-crafted inference attacks [51, 43, 17, 39, 61]. To mitigate privacy risks, differential privacy (DP) [14], a widely adopted standard for incorporating formal privacy guarantees, has been integrated into the FL algorithm [38]. In the context of FL, DP can be applied at two distinct protection levels: record-level DP, which protects individual data points within a client’s dataset, and client-level DP, which protects the participation of a client (i.e., the client’s entire dataset). This work focuses on achieving client-level differentially private FL (DPFL), as it typically yields better model utility than its record-level counterpart in cross-device settings [25]. In the literature, client-level DPFL is implemented using the Gaussian mechanism [14], where each client’s local model update is perturbed by adding calibrated Gaussian noise scaled according to a uniform privacy budget $\epsilon$ across all clients [5, 7, 50, 25, 29]. A smaller $\epsilon$ provides stronger privacy guarantees but requires injecting larger noise, which consequently leads to more severe model utility degradation. These perturbed model updates are typically aggregated using secure aggregation (e.g., [8]), which cryptographically ensures that the server can only access their sum without observing individual contributions. This dual protection yields a differentially private aggregated model update that prevents client-level privacy inference even with an adversarial server.
@@
However, in practice, clients often have heterogeneous privacy preferences, necessitating support for heterogeneous DP (HDP) [26, 33]. In the literature, Liu et al. [32] formally introduced the problem of FL with heterogeneous DP (HDPFL), where each client naturally has an individual privacy budget reflecting their privacy needs. In this setting, ensuring record-level HDP is relatively straightforward, and numerous studies have proposed to improve model utility [7, 33, 48, 58, 35, 36]. In contrast, client-level HDPFL remains under-explored. To achieve client-level DP with heterogeneous privacy requirements, conventional approaches such as DP-FedAvg [38] must satisfy the most stringent privacy budget among all clients, which severely limits overall model utility. A more practical alternative partitions clients into groups and enforces client-level DP at the group level. To improve the model utility in this scenario, recent efforts include manually adjusting per-group client sampling ratios [29], adjusting training rounds per group [11], and mitigating the influence of noisy per-group updates [32]. However, these approaches assume a fully trusted server, which is often unrealistic in settings that are vulnerable to privacy inference attacks. Moreover, they primarily rely on heuristic methods without rigorous theoretical analysis to optimize the privacy-utility trade-off.
@@
In this work, we aim to optimize the model utility in client-level HDPFL under a strong attack model where both the clients and the server are adversaries. We propose GDPFed, a novel client-level HDPFL approach in which clients are grouped by their privacy budgets, with client-level DP achieved at a group-wise level using each group’s minimum privacy budget rather than the global minimum. This design enables higher model utility while respecting heterogeneous privacy preferences. Building on this, we theoretically investigate how to maximize model utility in GDPFed while maintaining rigorous privacy guarantees. Through privacy and convergence analysis, we identify two key factors that influence convergence errors under fixed privacy budgets: (1) model dimensionality, as DP noise must be added to each model parameter, increasing total noise with model size; and (2) per-group client sampling ratios, which has a privacy amplification effect on the guarantees. To reduce dimensionality-induced noise, we incorporate model sparsification into GDPFed, which eliminates less significant model parameters for each group with minimal utility drop. We then optimize the per-group client sampling ratios towards minimizing the convergence error, which extends GDPFed to GDPFed+ with improved model utility. In summary, we make the following contributions:
@@
  • We propose GDPFed, a novel client-level DPFL algorithm for environments where both server and clients are honest-but-curious. GDPFed is specifically designed to improve model utility when clients have heterogeneous privacy preferences. By achieving client-level DP at a group-wise level, our approach mitigates privacy budget waste inherent in HDP settings, improving the model utility. GDPFed builds upon FedAvg framework, enabling seamless integration into existing FL systems.
@@
  • We conduct extensive empirical evaluations on multiple benchmark datasets of DPFL, thoroughly comparing our methods against state-of-the-art baseline methods. The results consistently demonstrate that GDPFed outperforms DP-FedAvg in HDP settings, while GDPFed+ further improves the model utility under the same privacy guarantee.
2 Preliminary and Related Work
@@
Attack Model. To achieve client-level DP, the literature typically assumes that the adversary is either honest-but-curious clients [7, 29, 38, 32] or, in a stronger setting, both the clients and the server [25, 50, 16, 27]. The adversary follows the prescribed FL protocol but remains curious about a target client’s private data and attempts to infer it from shared messages. In this work, we consider the latter, more challenging one.
@@
Federated Learning and FedAvg. In a typical FL system, a set of $n$ clients aim to collaboratively train a shared global model $\theta\in\mathbb{R}^{d}$ in an iterative manner under the coordination of a central server. Generally, the FL problem can be formulated as $\min_{\theta}(1/n)\sum_{i=1}^{n}f_{i}(\theta)$, where $f_{i}(\theta)=\mathbb{E}_{(z,y)\in D_{i}}l(\theta;z,y)$ represents the local learning objective of client $i$. Here, $l(\cdot)$ is the loss function, and $(z,y)$ is a datapoint sampled from the local dataset $D_{i}$ of client $i$. The classic method to solve the FL problem is known as Federated Averaging (FedAvg) [37]. Specifically, in each training round $t$, the server randomly selects a set of $r$ clients $\mathcal{S}^{t}$ with a client sampling ratio $q\in(0,1]$ without replacement to participate in the local training. Each client $i\in\mathcal{S}^{t}$ then downloads the latest global model $\theta^{t-1}$ from the server, refines the model for $\tau$ iterations towards optimizing its local objective to obtain an updated local model $\theta_{i}^{t}$ and then sends its local model updates $\Delta_{i}^{t}=\theta_{i}^{t}-\theta^{t-1}$ back to the server. The server refines the global model by averaging the local updates as $\theta^{t}=\theta^{t-1}+(1/r)\sum_{i\in\mathcal{S}^{t}}\Delta_{i}^{t}$. This process repeats for enough $T$ rounds to ensure that the global model converges. Since the server receives individual model updates from clients in each round, it poses a significant privacy risk, as a curious server can infer sensitive information from these updates.
@@
Differential Privacy. The DP mechanism [14, 41], especially the Gaussian mechanism (see the formal definition in Lemma 6), has been employed as a rigorous approach for mitigating privacy threats in FL [13, 25, 50]. We give the formal definition of classic $(\epsilon,\delta)$-DP in Definition 1.
@@
Given privacy budget $\epsilon>0$ and failure parameter $10\leq\delta<$ 1, a randomized mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if for any two adjacent datasets $D,D^{\prime}$, any subset of outputs ${O}\subseteq\text{range}(\mathcal{M})$ satisfies $\Pr[\mathcal{M}(D)\in{O}]\leq e^{\epsilon}\Pr[\mathcal{M}(D^{\prime})\in{O}]+\delta.$
@@
In this work, as we consider client-level DP, we define the adjacent datasets by adding or removing the entire local dataset of a client in FL. The privacy budget $\epsilon$ defines the upper bound on privacy loss in a DP mechanism. A smaller $\epsilon$ indicates stronger privacy protection but requires injecting more intense noise into the learning process, which can significantly impact model performance. Additionally, the failure parameter $\delta$ quantifies the probability that the DP guarantee may be violated. When $\delta=0$, the formulation $(\epsilon,\delta)$-DP simplifies to pure DP.
The standard $(\epsilon,\delta)$-DP provides a relatively loose composition bound, making it unsuitable for accurately tracking the cumulative privacy loss in complex iterative algorithms. Therefore, in this work, we adopt Rényi DP (RDP) [41], a relaxed variant of $(\epsilon,\delta)$-DP, to better quantify privacy loss over multiple rounds in DPFL. We provide the formal definition of RDP and its related properties used in this work in Appendix A.
@@
Client-level DP-FedAvg. Compared with record-level DP [2, 33, 34, 54], which aims to protect every individual record in a client’s dataset, client-level DP hides a single client’s overall contribution. To achieve client-level DP under our attack mdoel, one can use DP-FedAvg [38]: before transmitting the local model update $\Delta^{t}_{i}$ to the server at round $t$, each selected client clips its model update with a clipping threshold $C$, and adds small amount of DP noise drawn from $\mathcal{N}\left(0,C^{2}\sigma^{2}/r\cdot\mathbf{I}^{d}\right)$, where $\sigma^{2}$ is the noise multiplier. Notably, the noise multiplier $\sigma^{2}$ must be carefully calibrated to ensure that DP-FedAvg satisfies $(\epsilon,\delta)$-DP after $T$ training rounds. Theoretical analyses establish the relationship $\sigma^{2}=\Omega(q^{2}/\epsilon)$ [1, 42], implying that satisfying a smaller privacy budget $\epsilon$ necessitates injecting larger noise. Furthermore, DP-FedAvg benefits from privacy amplification via client subsampling [6], where in each client is independently selected with probability $q$ in every training round.
@@
After perturbing their updates locally, clients encrypt these noisy updates using a secure aggregation protocol (e.g., [8]) and send them to the server. Secure aggregation is a commonly used practice in client-level DPFL [25, 16, 27, 50], ensuring that a curious server only observes the aggregated sum of clients’ updates, without access to individual contributions. In this setting, the aggregated model update received by the server is already perturbed with Gaussian noise $\mathcal{N}\left(0,C^{2}\sigma^{2}\cdot\mathbf{I}^{d}\right)$. Finally, the global model is refined with the perturbed aggregated updates. If the server is assumed to be trusted [11, 29, 32], these model clipping and perturbation operations can be directly applied to the aggregated model update on the server side to prevent clients from inferring private information. We present the detailed DP-FedAvg algorithm in Algorithm 2 in Appendix B.3.
@@
The noise applied to model updates inherently reduces the utility of the global model. To mitigate this issue, numerous methods have been proposed, including model update regularization [5, 13, 50] to ensure more robust local updates, optimized client sampling [12, 52, 47] to select more informative clients, and sparsification [25, 13] to remove unnecessary noise. However, these methods consider a homogeneous DP setting, where all clients share the same privacy preference. In contrast, an HDP setting where clients have heterogeneous privacy preferences is more realistic and better aligned with practical deployment scenarios.
@@
FL with Heterogeneous Privacy Preferences. In practice, clients often have diverse privacy requirements due to varying policies or individual preferences, making it essential to consider FL under HDP [26]. Liu et al. [32] first formalized the problem of HDPFL, allowing each client to specify a unique privacy budget that reflects their preferences. In this setting, record-level HDP is straightforward to implement by calibrating the DP noise individually per client [7, 33, 48, 58, 32, 35]. For example, Boenisch et al. [7] proposed IDP-FedAvg, which assigns data sampling ratios and clipping thresholds based on each record’s privacy budget. However, achieving client-level HDP, where the goal is to protect a single client’s contribution from being inferred, poses greater challenges.
@@
Standard approaches such as DP-FedAvg [38] in this heterogeneous setting have to calibrate noise to satisfy the most stringent privacy requirement among clients, leading to excessive noise for clients with more relaxed privacy preferences and thus poor model utility [29]. A more privacy-efficient approach is to partition clients into groups based on their privacy budgets and ensure client-level DP within each group [29, 32, 11]. For instance, Kiani et al. [29] proposed a dynamic HDPFL framework where clients in different groups consume less privacy budget in early training rounds. Note that this method also proposes the formulation of client sampling ratio optimization, but does not address it in its design. Instead, they manually tune each group’s sampling ratio, limiting the method’s theoretical rigor. Another related method, Projected Federated Averaging (PFA) [32], retains updates from groups with high privacy budgets while projecting updates from low-budget groups onto the principal subspace learned from the high-budget group. Compared to PFA, our method improves the privacy-utility trade-off through both theoretical analysis and optimization techniques.
3 Federated Learning with Heterogeneous Group Client-Level DP

### v3

@@ Abstract
Federated Learning with client-level differential privacy (DP) provides a promising framework for collaboratively training models while rigorously protecting clients’ privacy. However, classic approaches like DP-FedAvg struggle when clients have heterogeneous privacy requirements, as they must uniformly enforce the strictest privacy level across all clients, leading to excessive DP noise and significant degradation in model utility. Existing methods to improve the model utility in such heterogeneous privacy settings often assume a trusted server and are largely heuristic, resulting in suboptimal performance and lacking strong theoretical foundations. In this work, we address these challenges under a practical attack model where both clients and the server are honest-but-curious. We propose GDPFed, which partitions clients into groups based on their privacy budgets and achieves client-level DP within each group to reduce the privacy budget waste and hence improve the model utility. Based on the privacy and convergence analysis of GDPFed, we find that the magnitude of DP noise depends on both model dimensionality and the per-group client sampling ratios. To further improve the performance of GDPFed, we introduce GDPFed+, which integrates model sparsification to eliminate unnecessary noise and optimizes per-group client sampling ratios to minimize convergence error. Extensive empirical evaluations on multiple benchmark datasets demonstrate the effectiveness of GDPFed+, showing substantial performance gains compared with state-of-the-art methods.
I Introduction
@@
Traditional centralized Machine Learning (ML) frameworks require collecting all training data at a single node (e.g., a central server), which raises serious privacy concerns and communication burdens. These issues are particularly evident in Internet of Things (IoT) systems, where massive numbers of resource-constrained devices continuously generate sensitive data and transmit it over bandwidth-limited networks. To address these challenges, Federated Learning (FL) [37] has emerged as a distributed ML paradigm that enables collaborative model training across decentralized data sources without centralizing raw data. FL has been widely applied in IoT-related domains such as healthcare [46] and remote sensing [34], where data privacy, communication efficiency, and on-device computation are critical. In a typical FL system, multiple local clients, such as IoT or edge devices, collaboratively train a shared global model under the coordination of a central server. In each training round, the server broadcasts the global model to a subset of clients, which update the model using their local data. The resulting model updates are then sent back to the server and aggregated to refine the global model. This iterative process continues until global model convergence.
@@
Although the FL paradigm keeps sensitive training data on clients, recent studies have shown that adversaries can still infer private information through well-crafted inference attacks [52, 43, 16, 39, 63]. To mitigate privacy risks, differential privacy (DP) [13], a widely adopted standard for incorporating formal privacy guarantees, has been integrated into the FL algorithm [38]. In the context of FL, DP can be applied at two distinct protection levels: record-level DP, which protects individual data points within a client’s dataset, and client-level DP, which protects the participation of a client (i.e., the client’s entire dataset). This work focuses on achieving client-level differentially private FL (DPFL), as it typically yields better model utility than its record-level counterpart in cross-device settings [25]. In the literature, client-level DPFL is typically implemented using the Gaussian mechanism [13], where each client’s model update is perturbed by adding Gaussian noise scaled according to a uniform privacy budget $\epsilon$ across all clients [4, 6, 51]. A smaller $\epsilon$ provides stronger privacy guarantees but requires injecting larger noise, which consequently leads to more severe model utility degradation. These perturbed model updates are typically aggregated using secure aggregation (e.g., [7]), which cryptographically ensures that the server can only access their sum without observing individual contributions. This dual protection yields a differentially private aggregated model update that prevents client-level privacy inference even with an adversarial server.
@@
However, in practice, clients often have heterogeneous privacy preferences, necessitating support for heterogeneous DP (HDP) [26]. In the literature, [32] formally introduced the problem of DPFL with heterogeneous privacy requirements for different clients (HDPFL), where each client naturally has an individual privacy budget reflecting their privacy needs. In this setting, ensuring record-level HDP is relatively straightforward, and numerous studies have proposed to improve model utility [6, 49, 59, 35, 36, 33]. In contrast, client-level HDPFL remains underexplored. To achieve client-level DP with heterogeneous privacy requirements, conventional approaches such as DP-FedAvg [38] must satisfy the most stringent privacy budget among all clients, which severely limits overall model utility. A more practical alternative partitions clients into groups and enforces client-level DP at the group level. To improve the model utility in this scenario, recent efforts include manually adjusting per-group client sampling ratios [29], adjusting per-group training rounds [10], and mitigating the influence of noisy per-group updates [32]. However, these approaches assume a fully trusted server, which is often unrealistic in settings that are vulnerable to privacy inference attacks. Moreover, they primarily rely on heuristic methods without rigorous theoretical analysis to optimize the privacy-utility trade-off.
@@
In this work, we aim to optimize the model utility in client-level HDPFL under a strong attack model where both the clients and the server are adversaries. We propose Group-based Differentially Private Federated Learning (GDPFed), a novel client-level HDPFL approach. In GDPFed, clients are grouped according to their privacy budgets, and client-level DP is enforced at the group level by using the minimum privacy budget within each group rather than the global minimum across all clients. This design preserves high model utility while accommodating heterogeneous privacy preferences across clients. Building on this, we theoretically investigate how to maximize model utility in GDPFed while maintaining rigorous privacy guarantees. Through privacy and convergence analyses, we identify two key factors influencing convergence errors under fixed privacy budgets: (1) the model dimensionality, since DP noise is injected into each model parameter, causing the total noise to scale with model size; and (2) the per-group client sampling ratio, which provides privacy amplification effects on the overall guarantees. To reduce dimensionality-induced noise, we incorporate model sparsification into GDPFed, which eliminates less significant model parameters for each group with minimal utility drop. We then optimize the per-group client sampling ratios towards minimizing the convergence error, which extends GDPFed to GDPFed+ with improved model utility. In summary, we make the following contributions:
@@
  • We propose GDPFed, a novel client-level HDPFL algorithm for environments where both server and clients are honest-but-curious. GDPFed is specifically designed to improve model utility when clients have heterogeneous privacy preferences. By achieving client-level DP at a group-wise level, our approach mitigates privacy budget waste inherent in HDP settings, improving the model utility. GDPFed builds upon FedAvg framework, enabling seamless integration into existing FL systems.
@@
  • We conduct extensive evaluations on multiple benchmark datasets of DPFL, thoroughly comparing our methods against state-of-the-art baseline methods. The results consistently demonstrate that GDPFed outperforms existing methods in HDP settings, while GDPFed+ further improves the model utility under the same privacy guarantee.
The remainder of this paper is organized as follows. Section II introduces the system settings and preliminaries on DP. Section III reviews related work. Section IV presents the proposed GDPFed method and its privacy analysis. Section V describes the enhanced GDPFed+ method and provides a rigorous convergence analysis. Section VI reports comprehensive experimental results. Finally, Section VII concludes the paper.
II System Settings and Preliminary
II-A Attack Model
@@
To achieve client-level DP, the literature typically assumes that the adversary is either honest-but-curious clients [6, 29, 38, 32] or, in a stronger setting, both the clients and the server [25, 51, 15, 27]. In this work, we consider the latter, more challenging one. Specifically, the adversary is assumed to follow the prescribed training protocol honestly but remains curious about the private data of a target client, attempting to infer it from the shared messages. In addition, certain clients may collude with the server or with one another to extract sensitive information about a specific victim client. Moreover, the adversary may also take the form of a passive external eavesdropper who can observe all shared messages during training but does not actively inject false messages or disrupt communication [25].
II-B Federated Learning and FedAvg
@@
In a typical FL system, a set of $n$ clients aim to collaboratively train a shared global model $\theta\in\mathbb{R}^{d}$ in an iterative manner under the coordination of a central server. Generally, the FL problem can be formulated as $\min_{\theta}(1/n)\sum_{i=1}^{n}f_{i}(\theta)$, where $f_{i}(\theta)=\mathbb{E}_{(z,y)\in D_{i}}l(\theta;z,y)$ represents the local learning objective of client $i$. Here, $l(\cdot)$ is the loss function, and $(z,y)$ is a datapoint sampled from the local dataset $D_{i}$ of client $i\in[n]$. The classic method to solve the FL problem is known as Federated Averaging (FedAvg) [37]. Specifically, in each training round $t$, the server randomly selects a set of $r$ clients $\mathcal{S}^{t}$ with a client sampling ratio $q\in(0,1]$ without replacement to participate in the local training. Each client $i\in\mathcal{S}^{t}$ then downloads the latest global model $\theta^{t-1}$ from the server, refines the model for $\tau$ iterations towards optimizing its local objective to obtain an updated local model $\theta_{i}^{t}$ and then sends its local model updates $\Delta_{i}^{t}=\theta_{i}^{t}-\theta^{t-1}$ back to the server. The server refines the global model by averaging the local updates as $\theta^{t}=\theta^{t-1}+(1/r)\sum_{i\in\mathcal{S}^{t}}\Delta_{i}^{t}$. This process repeats for enough $T$ rounds to ensure that the global model converges. Since the server receives individual model updates from clients in each round, it poses a significant privacy risk, as a curious server can infer sensitive information from these updates [43].
II-C Differential Privacy
@@
The DP mechanism [13, 42], especially the Gaussian mechanism (see the formal definition in Lemma 6 in Appendix A), has been employed as a rigorous approach for mitigating privacy threats in FL [12, 25, 51]. We give the formal definition of classic $(\epsilon,\delta)$-DP in Definition 1.
@@
Given privacy budget $\epsilon>0$ and failure parameter $0\leq\delta<1$, a randomized mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if for any two adjacent datasets $D,D^{\prime}$, any subset of outputs ${O}\subseteq\text{range}(\mathcal{M})$ satisfies $\Pr[\mathcal{M}(D)\in{O}]\leq e^{\epsilon}\Pr[\mathcal{M}(D^{\prime})\in{O}]+\delta.$
@@
In this work, as we consider client-level DP, we define the adjacent datasets by adding or removing the entire local dataset of a client in FL. The privacy budget $\epsilon$ defines the upper bound on privacy loss in a DP mechanism. A smaller $\epsilon$ indicates stronger privacy protection but requires injecting more intense noise into the learning process, which can significantly impact performance. Additionally, the failure parameter $\delta$ quantifies the probability that the DP guarantee may be violated.
III Related Work
III-A Client-level DP-FedAvg
@@
Compared with record-level DP [33, 55, 18, 62], which aims to protect every individual record in a client’s dataset, client-level DP hides a single client’s overall contribution. To achieve client-level DP under our attack model, one can use the DP-FedAvg algorithm [38], which is presented in Algorithm 3 in Appendix D. Specifically, before transmitting the local model update $\Delta^{t}_{i}$ to the server at round $t$, each selected client clips its model update with a clipping threshold $C$, and adds small amount of DP noise drawn from $\mathcal{N}\left(0,C^{2}\sigma^{2}/r\cdot\mathbf{I}^{d}\right)$, where $\sigma^{2}$ is the noise multiplier. Notably, the noise multiplier $\sigma^{2}$ must be carefully calibrated to ensure that DP-FedAvg satisfies $(\epsilon,\delta)$-DP after $T$ training rounds. Theoretical analyses have established the relationship $\sigma^{2}=\Omega(q^{2}/\epsilon)$ [1, 41], implying that satisfying a smaller $\epsilon$ necessitates injecting larger noise. Furthermore, DP-FedAvg benefits from privacy amplification via client subsampling [5], where each client is independently selected with probability $q$ in every training round.
@@
After perturbing their updates locally, clients encrypt these noisy updates using a secure aggregation protocol (e.g., [7]) and send them to the server. Secure aggregation is a commonly used practice in client-level DPFL [25, 15, 27, 51], ensuring that a curious server only observes the aggregated sum of clients’ updates, without access to individual contributions. In this setting, the aggregated model update received by the server is already perturbed with Gaussian noise $\mathcal{N}\left(0,C^{2}\sigma^{2}\cdot\mathbf{I}^{d}\right)$. Finally, the global model is refined with the perturbed aggregated updates. If the server is assumed to be trusted [10, 29, 32], these model clipping and perturbation operations can be directly applied to the aggregated model update on the server side to prevent clients from inferring private information.
@@
The noise applied to model updates inherently reduces the utility of the global model. To mitigate this issue, numerous methods have been proposed, including model update regularization [4, 12, 51] to ensure more robust local updates, optimized client sampling [11, 54, 48] to select more informative clients, and sparsification [25, 12, 53, 64] to remove unnecessary noise. For example, Wang et al. [53] sparsify each layer in model updates to remove unnecessary noise. However, these methods consider a homogeneous DP setting, where all clients share the same privacy preference. In contrast, an HDP setting where clients have heterogeneous privacy preferences is more realistic and better aligned with practical deployment scenarios.
III-B FL with Heterogeneous Privacy Preferences
@@
In practice, clients often have diverse privacy requirements due to varying policies or individual preferences, making it essential to consider FL under HDP [26]. Liu et al. [32] first formalized the problem of HDPFL, allowing each client to specify a unique privacy budget that reflects their preferences. In this setting, record-level HDP is straightforward to implement by calibrating the DP noise individually per client [6, 33, 49, 59, 32, 35]. For example, Boenisch et al. [6] proposed IDP-FedAvg, which assigns data sampling ratios and clipping thresholds based on each record’s privacy budget. Ma et al. [35] studied the client selection problem in FL with record-level HDP. However, achieving client-level HDP, where the goal is to protect a single client’s contribution from being inferred, poses greater challenges.
@@
Standard approaches such as DP-FedAvg [38] in this heterogeneous setting have to calibrate noise to satisfy the most stringent privacy requirement among clients, leading to excessive noise for clients with more relaxed privacy preferences and thus poor model utility [29]. A more privacy-efficient approach is to partition clients into groups based on their privacy budgets and ensure client-level DP within each group [29, 32, 10]. For instance, Kiani et al. [29] proposed a dynamic HDPFL framework where clients in different groups consume less privacy budget in early training rounds. While they formulate the client sampling ratio optimization problem, they manually tune each group’s sampling ratio, limiting the method’s theoretical rigor. Instead, our work is the first to provide a solution to the problem with theoretical justification. Another related method, Projected Federated Averaging (PFA) [32], retains updates from groups with high privacy budgets while projecting updates from low-budget groups onto the principal subspace learned from the high-budget group. Compared to PFA, our method improves the privacy-utility trade-off through both theoretical analysis and optimization techniques.
IV Federated Learning with Heterogeneous Group Client-Level DP
IV-A GDPFed